what is the difference quotient for f(x)=x(squared)-5x+8
Answer by stanbon(75887)

what is the difference quotient for
f(x)= x^2-5x+8
--------
f(x+h) = (x+h)^2 - 5(x+h) + 8
f(x+h) = x^2 + 2hx + h^2 - 5x - 5h + 8
------
f(x+h)-f(x) = 2hx +h^2 -5h
-----
[f(x+h)-f(x)]/h = 2x + h -5
=============================
Cheers,
Stan H.
============